# Algorithmic Fairness amid Social Determinants: Reflection, Characterization, and Approach

### v1

@@ 3.2 Quota-Based Admissions
Theorem 3.5 demonstrates how quota-based admissions can inadvertently disadvantage non-URM applicants from resource-poor regions, creating a new form of unfairness while attempting to address historical injustice. In particular, the larger the quota (larger $\eta_{\mathrm{quota}}$), the more spots are reserved for URM applicants (from both poor and rich regions), the more challenging for non-URM applicants in the poor region to be able to attend the selective college. In other words, non-URM applicants in the poor region, who face the same obstacles and disadvantages in contextual environments as their URM counterparts, are not reserved additional spots; on top of that, they have to compete with more advantaged peers (non-URM applicants from the rich region) over the spots that are already more limited. As we illustrate in Figure 1(a) Scenario (i), quota-based admission may result in a higher score threshold for non-URM in poor region than that for URM in rich region.
3.3 Holistic Review with Plus Factors
Putting aside the evolving jurisprudence [120, 121, 122], we aim to precisely characterize holistic review in terms of its implications on the distribution of benefits and burdens among individuals, when allocating the limited spots in selective college admissions. When taking into account of social determinants signified by Address Region, we show that holistic review with plus factors may benefit applicants from better-off areas more than those from less well-off areas:
Theorem 3.6 (Holistic Review with Plus Factors Benefits URM in Rich Region More).
Under Assumptions 3.13.4, let us denote with $\eta_{\dagger}<1$ the multiplicative coefficient on the scale parameter of Gamma distributions for URM applicants’ academic index scores, such that the perceived scores of URM applicants shift more probability density towards the high-score end. Let us denote with $q_{o}$ the default threshold for selective admission, and with $q_{\dagger}$ the threshold if the admission procedure is a holistic review with plus factors. Further assume that region-specific shape parameters satisfy $k^{(\mathrm{poor})}=k^{(\mathrm{rich})}=k_{o}$. Then, the increase in the probability of selective admission for URM applicants from the rich region, is larger than that for URM applicants from the poor region:
[…]
Theorem 3.6 characterizes different levels of benefits for URM applicants from different regions. Specifically, in terms of the increase in admission probability to the selective college, URM applicants from the rich region benefit more from the admission procedure that utilizes holistic review with plus factors, compared to URM applicants from the poor region. To better demonstrate our theoretical result, we provide illustrations in Figure 1(b).
As presented in top-row subfigures in Figure 1(b), at the original scale, the region-specific distributions of academic preparedness are the same for URM and non-URM applicants (Assumption 3.2). Holistic review with plus factors grants preference to URM applicants by perceiving their scores, at the distribution level, as if they were sampled from a distribution that is more concentrated at the high-score end (the plus-factor scale). Because of the limited availability in selective admissions, the threshold $q_{\dagger}$ for admission under holistic review with plus factors is more competitive than the default $q_{o}$, i.e., $q_{\dagger}<q_{o}$, for both URM and non-URM applicants. While non-URM applicants are assessed on the original scale, URM applicants are evaluated on a plus-factor scale. Under the Gamma parameterization (Assumption 3.3), this is equivalent to employing a more competitive threshold $q_{\dagger}$ for non-URM applicants but a less competitive one $q_{\dagger}/\eta_{\dagger}$ for URM applicants, where $q_{\dagger}<q_{o}<q_{\dagger}/\eta_{\dagger}$. Although the mathematical form of $q_{o}<\nicefrac{{k_{o}\ln(\theta^{(\mathrm{poor})}/\theta^{(\mathrm{rich})})}}{{(\nicefrac{{1}}{{\theta^{(\mathrm{rich})}}}-\nicefrac{{1}}{{\theta^{(\mathrm{poor})}}})}}$ appears convoluted, the condition itself is relatively mild. Graphically speaking, the spots at the selective college are limited such that the threshold $q_{o}$ does not reach the point where region-specific Gamma density curves (in the original scale) intersect, as depicted by $\widetilde{q}$ in Figure 1(b).
From the shaded areas in bottom-row subfigures in Figure 1(b), we can see that the increased admission probability for URM groups comes with a corresponding reduction in that for non-URM groups. However, such redistribution benefits URM applicants in the rich region more than those in the poor region, essentially disadvantaging URM applicants in less well-off areas.
3.4 Top-Percentage Plans
Taking into account the demographic composition of applicants and the number of available spots at the selective college, we characterize the difference between top-percentage plans compared to the default selective admission. When explicitly considering the role of Address Region in applicants’ academic preparedness, we show that the redistribution of limited selective admissions, as implied by top-percentage plans, is carried out by reallocating availability from the rich region to the poor region, regardless of the demographic group of applicants:
Theorem 3.7 (Top-Percentage Plans Reallocate Spots from Rich Region to Poor Region).
Under Assumptions 3.13.4, let us denote with $q_{o}$ the default threshold for selective admission, and with $q^{(\mathrm{poor})}$ and $q^{(\mathrm{rich})}$ the thresholds for poor and rich regions, respectively, if top-percentage plans are employed. Then, the increase in selective admissions (in terms of counts) for applicants from the poor region, comes from spots reallocated out of the rich region. This redistribution is a result of the top-percentage plans, and is not relevant to applicants’ demographic group:
$\small\begin{split}\big{(}n_{a}^{(\mathrm{poor})}+n_{a^{\prime}}^{(\mathrm{poor})}\big{)}\big{[}F^{(\mathrm{poor})}(q^{(\mathrm{poor})})-F^{(\mathrm{poor})}(q^{(o)})\big{]}=\big{(}n_{a}^{(\mathrm{rich})}+n_{a^{\prime}}^{(\mathrm{rich})}\big{)}\big{[}F^{(\mathrm{rich})}(q^{(o)})-F^{(\mathrm{rich})}(q^{(\mathrm{rich})})\big{]}.\end{split}$
Theorem 3.7 characterizes the reallocation of the selective admission spots performed by top-percentage plans. In Figure 1(c), we use shaded areas to illustrate the transfer of admission opportunity (in terms of the region-wise probability of selective admission) from the rich region to the poor region. The additional selective admissions gained by the poor region, compared to the default setting, are distributed proportionally to the natural demographic composition of each group.
4 Experiments
Commonly used datasets and benchmarks in algorithmic fairness literature tend to omit variables related to social determinants (as we discussed in Section 2.3.3). However, the relative absence of comprehensive measurements does not render our framework unnecessary or ineffective. In this section, we demonstrate how to apply our analytical framework using the information available. We consider the publicly-available statistics for freshmen admissions to University of California (UC), and reason about underlying academic preparedness from potential regions.
4.1 Formulation of the Optimization Problem
Due to legal and ethical reasons, the released data only contains summary statistics, and the detailed application or admission data is not publicly available. Nevertheless, we aim to utilize the information available and estimate region-specific academic preparedness.
We do not regard race as a determinant of academic preparedness (Assumption 3.2), and incorporate the Gamma parameterization for region-specific distribution of academic preparedness among applicants (Assumption 3.3). Both the number of regions and demographic groups can take on values beyond the binary case. After specifying the number of regions, we formulate a constrained optimization problem to solve for region-specific shape and scale parameters, as well as demographic compositions across regions.555 The data is obtained from UC undergrad admissions summary and freshmen fall admissions summary. We provide data descriptions, formulation of constrained optimization, as well as additional analyses in Appendix C.
(a) Score densities
(b) Race- and region- specific composition
(c) Quantile-specific composition
Figure 2: Visualization of constrained optimization results fitted on University of California application and admission summary statistics. Panel (a): region-specific and overall densities of academic preparedness. Panel (b): for each group, the region-specific compositions of application and admission proportions (left four subplots); for each region, the demographic composition of applicants (right subplot). Panel (c): for specific quantiles of interest, the region composition of admitted students (in terms of the proportion among all applicants).
4.2 Experimental Results
Because of the lack of individual-level data, the optimization problem can remain under-constrained due to the limited information available provided by summary statistics. In practice, we solve the constrained optimization problem to match the estimation with the university-wide statistics of capped and weighted high-school GPA scores (from year 2023).666 Our implementation can be found at the Github repository https://github.com/zeyutang/FairnessAmidSocialDeterminants. We consider demographic groups recorded in the data, and limit the number of potential regions to three to avoid overfitting of summary statistics. In Figure 2, we present visualizations of the result of the constrained optimization, including the estimated region-wise and demographic composition of applicants, $n_{a}^{(r)}$, the parameters in region-specific Gamma distributions, $k^{(r)}$ and $\theta^{(r)}$, and the corresponding score thresholds $q^{(r)}$, where region $r\in\{0,1,2\}$ and race $a\in\{\text{African American},\text{Asian},\text{Hispanic/Latino(a)},\text{White},\text{Others}\}$. We acknowledge the limitations of our constrained optimization approach given the aggregated nature of the available data. Individual-level data would enable more robust validation of our framework’s applicability.777 In Appendix D, we demonstrate how address information can be used to link existing datasets to socioeconomic status indicators, potentially enabling the development of more context-aware fairness benchmarks.
In Figure 2(a), we present region-specific densities of academic preparedness, as well as the overall density if we consider all applicants. The distinct shapes of region-specific densities reflect the varying influences on applicants’ academic preparedness across different regions. For instance, the densities of Region 0 (blue) and Region 2 (green) concentrate more at the high-score end, compared to Region 1 (orange), indicating the more positive influence on applicant’s academic preparedness. In Figure 2(b), in the left four subplots, for different demographic groups we present region-specific compositions of application and admission proportions; in the right-hand-side subplot, we present the demographic composition of applicants within each region. In Figure 2(c), we present the proportion (among all applicants) of admitted students whose scores are above specific quantiles. As we can see from Figure 2, there is a correlation between race and social determinants, as indicated by different academic preparedness across regions, and also by the disproportionate demographic compositions of admission even if the procedure does not utilize race (as per 1996 California Proposition 209).
5 Concluding Remarks
Algorithmic fairness research has largely focused on sensitive attributes, leaving important roles of social determinants under-explored. In this paper, we address this gap by introducing formal and quantitative rigor into a space shaped primarily by qualitative insights. Using college admissions as a concrete setting, we model region as a surrogate for social determinants and apply Gamma distribution parameterization to capture the effects of potential structural injustice. Despite its simplicity, our approach recovers nuanced findings aligned with prior qualitative debates, which previous quantitative approaches are not able to produce. Our results suggest that fairness mitigation strategies based solely on sensitive attributes risk introducing or reinforcing structural injustice.
Because social determinants correlate with sensitive attributes, explicitly considering social determinants through which structural injustice potentially perpetuates can help us better understand the underlying data generating process. This, in turn, facilitates more precise and comprehensive fairness characterization and mitigation strategies. Incorporating social determinants also makes it more transparent to see benefits and burdens experienced by individuals with different demographic backgrounds and contextual environments, when they are subjected to different algorithmic decision-making procedures. Policymakers should consider mandating the collection and analysis of social determinants alongside sensitive attributes when evaluating algorithmic systems for fairness compliance.
This framework could be similarly applied to other algorithmic decision-making contexts such as healthcare resource allocation, lending decisions, and hiring processes, where social determinants also play crucial roles. Future works naturally include designing and utilizing appropriate measurements of social determinants to develop fairness auditing and mitigation strategies, so that we can achieve fairness in an effective, principled, and transparent way.
Supplement to
“Algorithmic Fairness amid Social Determinants:
Reflection, Characterization, and Approach”
[…]
$\printcontents[supp]$l1
Table of Contents: Appendix

### v2

@@ 4.2 Structural Injustice Implications of A Classic Mitigation Strategy
Specifically, as the quota increases ($\eta_{\mathrm{quota}}$ grows), more seats are reserved for URM applicants (from both poor and rich regions), the more challenging for non-URM applicants in the poor region to be able to attend the selective college. Non-URM applicants in the poor region, who face the same obstacles and disadvantages in contextual environments as their URM counterparts, are not reserved additional spots; on top of that, they have to compete with more advantaged peers (non-URM applicants from the rich region) for a shrinking pool of openings.
More importantly, and counter-intuitively, this unintended harm is less likely when existing structural injustice is more severe, as captured by a larger degree of stochastic dominance on the left-hand side of Equation (1). However, as structural justice improves and regional dominance diminishes, quota-based mitigation becomes increasingly likely to generate new forms of structural injustice, since the inequality in Equation (1) can get violated. Moreover, the right-hand side of Equation (1) grows with the quota size $\eta_{\mathrm{quota}}$, indicating that stronger sensitive-attribute-centered intervention amplifies this unintended harm more drastically.
Unlike traditional arguments against affirmative action that seek to ignore certain sensitive attributes, we argue that ignoring the interaction between sensitive attributes and social determinants harms the most disadvantaged subsets of all demographic groups.
Takeaway: There is a paradoxical relationship between sensitive-attribute-centered mitigation and structural injustice: certain mitigation strategies are less likely to introduce new disadvantage when existing structural injustice is more severe, but increasingly generate new forms of structural injustice as conditions improve.
(a) PUMAs w/ low-level area deprivation
(b) PUMAs w/ mid-level area deprivation
(c) PUMAs w/ high-level area deprivation
Figure 1: Histogram of annual income for African American women residing in different areas. Sensitive attributes (race and sex) are shared across the subfigures, whereas social determinants are not (higher ADI indicates higher area deprivation).
(a) Rich region
(b) Poor region
(c) Semi-synthetic screening settings
(d) Quantitative comparison of scenarios
Figure 2: Panels (a) and (b): Histogram of the age at the first-ever breast cancer screening for White women residing in different areas. Sensitive attributes (race and sex) are shared, whereas social determinants are not: the rich region corresponds to ADI $\in[0,25)$ and the poor region corresponds to ADI $\in[75,100)$. Panels (c) and (d): We conduct a semi-synthetic simulation of breast cancer onset and screening to evaluate how policy interventions targeting social determinants affect early detection outcomes. Panel (c) shows that adopting improved (rich-region) screening pattern in the poor region shifts detections earlier. Panel (d) demonstrates quantitatively that such structural improvements yield non-trivial gains in early detections across allocation scenarios (with a same total number of screenings).
5 Empirical Analyses
In this section, we provide empirical analyses of real-world interplay between sensitive attributes and social determinants. In Section 5.1, we demonstrate the insufficiency of (intersectional) sensitive attributes when capturing disadvantage on the U.S. Census data (Census Bureau, 2009, 2014, 2022). In Section 5.2, we present a semi-synthetic analysis in a high-stakes domain application regarding breast cancer screening within an integrated U.S. healthcare system.
5.1 Enhancing Census Data Product Though Linking Socio-Economic Status Indices
We retrieve the public use microdata sample (PUMS) data from the U.S. Census Bureau (Census Bureau, 2023), and provide visualizations of different Public Use Microdata Areas (PUMAs) in California based on the 2023 U.S. Census PUMS data. Each PUMA contains at least $100,000$ residents and provides reliable, detailed demographic, economic, and housing statistics at a sub-state level while also protecting the confidentiality of respondents. In order to link PUMAs to indicators of (area-level) social determinants, we also retrieve the updated Area Deprivation Index (ADI). [^3] https://www.neighborhoodatlas.medicine.wisc.edu/
In Figure 1, we present the histogram of annual income for African American women residing in PUMAs with different ADI levels. As we can see, although the demographic information reflects the intersectional characteristics of individuals (race and sex), the social determinants in different regions still play a nontrivial role in shaping the income distribution. For instance, in PUMAs with higher ADI levels, the income distribution is more skewed towards lower income levels, indicating that individuals in these areas may face more significant economic challenges compared to those in PUMAs with lower ADI levels. We provide further analyses on PUMAs in Appendix C.
5.2 Semi-Synthetic Analysis in a High-Stakes Domain: Breast Cancer Care
In addition to the demographic analyses based on the U.S. census data, we also analyze de-identified patient records from OSF HealthCare, an integrated healthcare system spanning multiple hospitals and outpatient facilities across a multi-county region, serving both urban and rural communities.Focusing on breast cancer screening and care from 2012–2022, we linked cancer registry-derived mammography logistics data (e.g., screening and diagnostic mammography dates and locations) with demographic and clinical variables (e.g., birth date, gender, race/ethnicity, residential ZIP code, residential area socio-economic index values) using anonymized patient identifiers to ensure full de-identification. The final dataset contains approximately 54,000 screening mammography records from around 45,000 unique patients.
In Figures 2(a) and 2(b), we present the distribution of age at first breast cancer screening for White women across regions within the healthcare system. As we can see, even among individuals with identical sensitive attributes (White women), the age at first-ever breast cancer screening differs by several years (e.g., over 3 years in the mean and nearly 5 years in the median). Because the integrated healthcare system applies a uniform screening guideline across regions, these gaps cannot be attributed to region-specific recommendations. [^4] Nevertheless, across the United States, breast cancer screening guidelines broadly agree that screening mammography should begin around age 40 for average-risk individuals, but differ on recommended frequency. Instead, they are consistent with structural disadvantages, such as differences in transportation infrastructure, access to care, or other contextual barriers, that shape screening uptake. The fact that the guideline is not explicitly social-determinant-aware indicates the inadequacy of fairness through unawareness, mirroring prior insights regarding sensitive attributes (Dwork et al., 2012; Lipton et al., 2018), now situated in the context of social determinants.
To illustrate the impact of policy interventions targeting social determinants, we perform a semi-synthetic simulation. We use 100,000 particles (the minimum PUMA population size) per region (rich and poor) with ages uniformly distributed from 20-99. For each particle in both regions, we simulate whether and when breast cancer develops by sampling year-by-year from age 20 onwards using SEER age-specific incidence rates (2018-2022) [^5] https://seer.cancer.gov/statistics-network/explorer/application.html?site=55&data_type=1&graph_type=3&compareBy=race&chk_race_1=1&rate_type=2&sex=3&advopt_precision=1&advopt_show_ci=on , so that the simulated age-specific onset rate is consistent with epidemiological data. We then allocate 10,000 screening slots across regions according to different policy scenarios. For each screening, we sample a “first-ever screening age” from empirical distributions derived from real data. An “early detection” benefit occurs when a particle is screened at or before her cancer onset age (assuming reliable periodic screenings after the initial screening), representing cases where screening enables earlier breast cancer care.
Figure 2(c) shows screening ages at which early detections occur. The “existing screening pattern” follows the poor region’s empirical distribution of age at first-ever screening (Figure 2(b)), and the “improved screening pattern” follows that of the rich region (Figure 2(a)). Both screening patterns allocate all 10,000 screenings to the poor region, differing only in the first-screening age distribution used. From the overlaid histograms, we can see that improvements in social determinants, instantiated through improved screening patterns in the poor region, lead to earlier detection timing.
Figure 2(d) compares four allocation scenarios for 10,000 new screenings (500 simulation runs): equal split (5,000 each region) versus all to the poor region, crossed with existing versus improved screening patterns. Shifting to an improved screening pattern yields a nontrivial gain, for example, increasing early detections from $1,367\pm 33$ to $1,461\pm 36$ per 10,000 screenings allocated to the poor region. [^6] We acknowledge that our semi-synthetic simulation abstracts from the full complexity of underlying epidemiological processes. Beyond first-time screening, incidence is also shaped by additional factors, e.g., genetic (Collins & Politopoulos, 2011; Shiovitz & Korde, 2015; Pal et al., 2024), epigenetic (Kim et al., 2023; Prabhu et al., 2024), that may shift observed onset.
Takeaway: We empirically demonstrate the instantiation of social determinants beyond sensitive attributes and, via semi-synthetic experiments, show the nontrivial benefits of policies that improve them.
6 Alternative Views
6.1 Viewpoint: Social Determinants Are “Just Another” Sensitive Attribute from a Technical Perspective
This alternative view states that, technically, social determinants can be treated as “just another” sensitive attribute, and therefore, there is no need to capture structural injustice specifically through social determinants:
Under associative fairness metrics, the influence of social determinants is instantiated as distribution shifts across domains, and can be addressed through domain adaptation mechanisms.
We respond to this alternative view by arguing that social determinants are not inherently individual-level attributes that merely shift across domains. Rather, they reflect structural injustices that shape individuals opportunities and outcomes in ways that are detached from specific identities or demographic group memberships. Domain adaptation and transfer learning style fairness approaches typically ask (Madras et al., 2018; Coston et al., 2019; Schumann et al., 2019; Creager et al., 2021; Mukherjee et al., 2022; Teo et al., 2023): “How can we learn predictors that eliminate empirical measures of discrimination under distribution shifts over observed individual attributes, so that disparities across individuals or demographic groups is robustly mitigated?” This framing treats heterogeneity across domains as variations in observable distributions, while keeping discrimination defined relative to individual/group identity.
In contrast, inquiries centered on social determinants ask a fundamentally different question: “How can we identify and quantify latent structural factors that give rise to these domains in the first place, i.e., factors whose (dis)advantageous effects on individuals are not tightly bound to individual identities or demographic categories?” Essentially, the goal is to capture both within-group heterogeneity (different levels of advantage among individuals of the same demographic group) and cross-group structural burdens (similar disadvantages experienced across different demographic groups). Social determinants operate at a structural level that is not intended to be eliminated through distributional robustness of a predictor, and thus cannot be treated as merely another sensitive attribute. It is also important to distinguish our position from redlining (Rothstein, 2017): rather than using correlated non-sensitive attributes (e.g., social determinants) as proxies for sensitive attributes to facilitate discrimination, we seek to audit how structural injustice is instantiated so it can be systematically addressed.
6.2 Viewpoint: Causal Effects of Sensitive Attributes Already Encompass Social Determinants
This alternative view states that, the influence of social determinants are captured either implicitly or explicitly by the causal effects of sensitive attributes: For causal fairness metrics, social determinants are modeled either implicitly (encoded in directed edges from sensitive attributes to downstream variables) or explicitly (as mediators along the causal pathways originating from sensitive attributes), and therefore, their influences are encompassed in causal effects of sensitive attributes.
We respond by arguing that causal effects of sensitive attributes do not correspond to interventions on the underlying structural conditions, and therefore, cannot recover the causal influence from social determinants.
First, for the “implicit” case, representing social determinants through directed causal edges from sensitive attributes to downstream variables creates a mismatch between the technical meaning of the causal model and the fairness objectives it is intended to capture. Let us consider $\texttt{Race}\rightarrow\texttt{Education Status}$ (Kusner et al., 2017; Nabi & Shpitser, 2018; Chiappa, 2019; Nabi et al., 2022) as an example. Even if we treat directed edges only as reductive summaries (e.g., encoding the influence of unobserved factors such as neighborhood affluence, geographic location, or institutional access) instead of asserting direct causal effect, there is a mismatch: while the edge stands in for latent contextual influences, the contrast operates on sensitive attributes themselves rather than on the underlying social determinants it proxies.
Second, for the “explicit” case, while previous path-specific fairness approaches (Zhang & Bareinboim, 2018a; Chiappa, 2019; Wu et al., 2019) can potentially include social-determinant variables as mediators along causal pathways from sensitive attributes to outcomes, this framing treats such variables as static downstream features rather than as shifting structural levers that could be the target of intervention. Moreover, social determinants need not be downstream variables of sensitive attributes. For example, race or sex is not a plausible ancestor node of environmental or contextual variables, which affect individuals in ways that are not mediated by, nor specific to, demographic memberships.
Takeaway: Both alternative views collapse social determinants into sensitive attributes, but doing so obscures their fundamentally non-interchangeable conceptual roles and leads to misaligned technical abstractions.
7 Call to Actions and Concluding Remarks
Addressing structural injustice in ML fairness demands coordinated advances in data governance, dynamic metric design, and explicit characterization of underlying data generating processes. In this section, we outline these actionable pillars.
Pillar 1 Data governance: Privacy-Aware Use of Social Determinant Data
Institutions should establish tiered data governance frameworks that distinguish between raw identifiers, privacy-preserving aggregates, and analytically derived contextual features. Concretely, raw identifiers (e.g., exact address/community, census tract) should be stored in restricted data locations with audited access, while model development relies on privacy-preserving contextual variables such as neighborhood deprivation indices, environmental exposure measures, or resource accessibility scores computed via differential privacy (Dwork et al., 2014) mechanisms. Data curators should document how social determinant variables are constructed, quantify privacy risk, and provide standardized contextual feature sets that can be safely integrated into datasets and benchmarks.
Pillar 2 Metric Design: Dynamic and Context-Adaptable Fairness Metrics
Fairness metrics should be redesigned to track disparities across evolving structural conditions rather than static demographic groupings or individual identities alone. Practically, this involves coupling individual-level attributes with time-varying contextual indicators (e.g., level of environmental pollution, policy exposure, access to care) and computing disparity measures conditional on these structural states. For example, as a counterpart to Demographic Parity (Calders et al., 2009; Dwork et al., 2012; Zemel et al., 2013; Feldman et al., 2015), Social Determinant Parity could evaluate (predicted) outcome disparities across area deprivation indices, infrastructure access levels, or regions with differing policy exposure. In addition, longitudinal formulations of Social Determinant Parity can track how outcome disparities change as contextual variables shift over time, enabling ongoing diagnostic of how algorithmic systems interact with changing social environments.
Pillar 3 Underlying Data Generating Processes: Identifying Intervention Levers
Causal fairness frameworks should treat social determinants, especially community- and context-level conditions, as explicit intervention targets. Concretely, causal fairness approaches should construct multi-layer causal models in which individual outcomes are jointly influenced by personal attributes and contextual variables such as neighborhood resources, environmental exposures, and policy exposures. This shift also necessitates leveraging causal representation learning (Xie et al., 2020; Schölkopf et al., 2021; Zheng & Zhang, 2023; Xie et al., 2024; Ng et al., 2024; Zhang et al., 2024; Dai et al., 2025) to actively seek out latent factors that could reflect structural characteristics, as well as hybrid inference methods that jointly perform causal inference (Pearl, 2009; Peters et al., 2017; Hernán & Robins, 2020) and relational inference between individuals and their surrounding contexts (Friedman et al., 1999; Maier et al., 2010; Ahsan et al., 2023).
Conclusion: Addressing structural injustice requires ML fairness research to provide technical tools that extend beyond addressing sensitive attributes, enabling the identification of effective structural levers and the rigorous measurement of progress toward structural justice, so that policy interventions can be better informed and evaluated.